OpenStudy (anonymous):

Solve the inequality. |r + 3| ≥ 7

OpenStudy (anonymous):

My favourite... for values of r + 3 ≥ 7, |r+3| = r+3 for values of r + 3 < 7, |r+3| = -(r+3) = -r-3 so for values of r ≥ 4, solve for r + 3 ≥ 7 so r ≥ 4. That is all values greater than equal to 4 for values of r < 4, solve for - r - 3 ≥ 7 -r ≥ 10 r ≤ -10 so all values less than or equal to -10 so to sum it up, the answer is r ≤ -10, and r ≥ 4

OpenStudy (zepp):

@psujono, if r = -9, the inequalify would still work..

OpenStudy (zepp):

Oh yeah, nvm